# Increasing Trust in Language Models through the Reuse of Verified Circuits

### v8

@@ Appendix B Appendix: Model Configuration
Addition, subtraction and mixed (addition and subtraction) training experiments were done in a Colab notebook. The Colab runs on a T4 GPU. Training takes up to 35 mins. The key parameters (and their common configurations) are:
@@
Each digit is represented as a separate token. (Liu and Low, 2023) state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers". The model’s vocabulary contains 14 tokens ( 0, .., 9, +, -, =, *, / ) to enable this and planned future investigations.
@@ Appendix C Appendix: Model Loss
In our experimental models, the number of digits in the question varies from 5 to 10, the number of layers varies from 1 to 4, the number of heads varies from 3 to 4. Each experimental model’s loss is detailed in Tab. 6.
@@ Appendix E Appendix: Experimental models
Twenty-one experiment models were trained and analyzed (refer Tab. 6). The models and output of analysis are available on HuggingFace at https://huggingface.co/PhilipQuirke/VerifiedArithmetic
@@
Table 6: Main experimental models studied. The number of addition and subtraction failures per million questions is shown. The best 5-, 6- and 10-digit models are bolded.
For each model the ’VerifiedArithmeticTrain’ Colab notebook generates two files:
  • A “XXXXXX.pth" file containing the model weights
  • A “XXXXXX_train.json" file containing configuration information and training loss data
While, for each model the ’VerifiedArithmeticAnalysis’ Colab notebook generates two more files:
  • A “XXXXXX_behavior.json" file containing generic “behavior" facts learnt about the model by the Colab e.g. P18L0H0 attends to tokens D3 and D’3
  • A “XXXXXX_maths.json" file containing “maths-specific" facts learnt about the model by the Colab e.g. P18L0H0 performs the A3.SC sub-task.
@@ Appendix G Appendix: Complexity
Table 8: We categorise addition questions into non-overlapping “calculation complexity" quanta, ordered by increased computational difficulty (and decreasing occurrence frequency). Five-digit addition questions quanta are S0 to S5. Ten-digit addition question quanta are S0 to S10. S10’s frequency is $\sim 3e-4$ showing the need to enrich training data for rare edge cases.
@@ Appendix H Appendix: Addition Hypothesis 1
  • The long double staircase still finalises each answer digit’s calculation.
@@ Appendix I Appendix: Addition Hypothesis 2
Table 10: Converting a $A_{n}$.T value into a SA , SC or SS value.
Our notation shorthand for one-digit-accuracy these “conversion" bigram mappings is:
@@
  • The calculation of A1.ST2 in P14.L0.H1 is a interesting case. The model needs A1.ST2 for A2 accuracy. The model could simply reuse the accurate A1.ST2 value calculated in P10. Activation patching shows that it does not. Instead the P14 attention heads calculate A1.ST1 from D1 and D’1 directly, and only relies on the P10.D1.ST2 value in the case where A1.ST2 != A1.ST. That is, the calculation is “use P14.A1.ST1 value else use A1.ST2 values". This aligns with the model learning the P10.A1.ST calculation early in training (for 90% accuracy) and later learning that P10.A1.ST2 contains additional information it can use to get to six nines accuracy.
@@
To test the hypothesis 3 mapping of the mathematical framework (casual abstraction) to the model attention heads, various “interchange interventions" (aka activation patching) experiments were performed on the model, where
@@
Table 13: We define “positive-answer subtraction" and “negative-answer subtraction" calculation complexity quanta that parallel the addition quanta.
@@
We experimented with three approaches to re-using the trained addition model in the “mixed" (addition and subtraction) model:
@@
  • Freeze Attention: As per “Initialize Only", but also every 100 training steps recopy the attention head weights from the addition model into the partially-trained mixed model.
  • Freeze All: As per “Initialize Only", but also every 100 training steps recopy the entire addition model (attention heads and MLP layers) into the partially-trained mixed model.
Our intuition was that “Initialize Only" would give the mixed model the most freedom to learn new algorithms, but that the “Freeze Attention" and “Freeze All" approaches would make the resulting trained mixed model easier to interpret (as we could reuse our addition model insights).
After experimentation we found that the “Initialize Only" approach was the only one that quickly trained to be able to do both addition and subtraction accurately. We concluded that the other two methods constrainws the model’s ability to learn new algorithms too much.
We also experimented with “where" in the model we inserted the addition (6-digit, 2-layer, 3-head) model into the slightly larger (6-digit, 3-layer, 4-head) mixed model. That is, do we initialize the first 2 layers or the last 2 layers of the mixed model? Also do we initialize the first 3 attention heads or the last 3 attention heads of the mixed model? Our intuition was that initializing the first layers and heads would be more likely to cause the model to re-use the addition circuits adding interpretability, so we used this approach.
@@
  • The model contains a new sub-task that stands out: The algorithm relies on calculations done at token position P0, when the model has only seen one question token! What information can the model gather from just the first token? Intuitively, if the first token is a “8" or “9" then the first answer token is more likely to be a “+" (and not a “-"). The model uses this heuristic even though this probabilistic information is sometimes incorrect and so will work against the model achieving very low loss.

### v9

@@ Appendix B Appendix: Model Configuration
Addition, subtraction and mixed (addition and subtraction) training experiments were done in a Colab notebook. The Colab runs on a T4 GPU. Each training run takes up to 60 mins. The key parameters (and their common configurations) are:
@@
Each digit is represented as a separate token. (Liu & Low, 2023) state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers”. The model’s vocabulary contains 14 tokens ( 0, .., 9, +, -, =, *, / ) to enable this and planned future investigations.
@@ Appendix C Appendix: Model Loss
In our experimental models, the number of digits in the question varies from 5 to 13, the number of layers varies from 1 to 4, the number of heads varies from 3 to 4. Each experimental model’s loss is detailed in Tabs. 6 and  7.
@@ Appendix E Appendix: Experimental models
Forty-four models were trained and analyzed (Refer Tabs. 6 and  7). The models and output of analysis are available on HuggingFace at https://huggingface.co/PhilipQuirke in the quanta_maths_* collections to support further research in AI Safety.
Table 6: Addition-only models studied. The number of addition failures per million questions is shown. The number of useful attention heads at token position and useful MLP layers at token position are shown - summarizing the data shown in figures like Fig 7.
Num Digits Num Layers Num Heads Train Steps Train Seed Train loss Addition Fails / M Heads used MLPs used
5 1 3 30K 372001 9.4e-2 12621 15 6
5 2 3 15K 372001 1.6e-8 0 30 16
5 2 3 40K 372001 2.0e-9 0 22 15
6 2 3 15K 372001 1.7e-8 2 31 17
6 2 3 20K 173289 1.5e-8 0 28 17
6 2 3 20K 572091 7.0e-9 0 35 17
6 2 3 40K 372001 2.0e-9 0 29 17
7 2 3 45K 173289 3.0e-9 0 31 20
8 2 3 45K 173289 3.0e-9 0 35 22
9 2 3 45K 173289 3.0e-9 0 54 27
10 2 3 40K 572091 7.0e-9 0 44 28
11 2 3 50K 173289 8.0e-9 2 56 29
12 2 3 50K 173289 5.0e-9 3 50 33
13 2 3 50K 173289 6.3e-8 1 66 31
14 2 3 60K 173289 5.5e-6 199 68 35
15 2 3 80K 572091 8.6e-8 10 93 58
Table 7: Subtraction-only and mixed models studied. The number of addition and subtraction failures per million questions is shown. The number of useful attention heads at token position and useful MLP layers at token position are shown - summarizing the data shown in figures like Fig 7.
@@
For each model the ’QuantaMathsTrain’ Colab notebook generates two files:
  • A “model.pth” file containing the model weights
  • A “training_loss.json” file containing configuration information and training loss data
While, for each model the ’QuantaMathsAnalysis’ Colab notebook generates two more files:
  • A “behaviors.json” file containing generic “behavior” facts learnt about the model by the Colab e.g. P18L0H0 attends to tokens D3 and D’3
  • A “features.json” file containing maths-specific ”feature” facts learnt about the model by the Colab e.g. P18L0H0 performs the A3.SC subtask.
@@ Appendix G Appendix: Complexity
Table 9: We categorise addition questions into non-overlapping “calculation complexity” quanta, ordered by increased computational difficulty (and decreasing occurrence frequency). Five-digit addition questions quanta are S0 to S5. Ten-digit addition question quanta are S0 to S10. S10’s frequency is $\sim 3e-4$ showing the need to enrich training data for rare edge cases.
@@ Appendix H Appendix: Addition Hypothesis 1
  • The long double staircase still finalizes each answer digit’s calculation.
@@ Appendix I Appendix: Addition Hypothesis 2
Our notation shorthand for one-digit-accuracy these “conversion” bigram mappings is:
@@
  • The calculation of A1.ST2 in P14.L0.H1 is a interesting case. The model needs A1.ST2 for A2 accuracy. The model could simply reuse the accurate A1.ST2 value calculated in P10. Activation patching shows that it does not. Instead the P14 attention heads calculate A1.ST1 from D1 and D’1 directly, and only relies on the P10.D1.ST2 value in the case where A1.ST2 != A1.ST. That is, the calculation is “use P14.A1.ST1 value else use A1.ST2 values”. This aligns with the model learning the P10.A1.ST calculation early in training (for 90% accuracy) and later learning that P10.A1.ST2 contains additional information it can use to get to >99.999% accuracy.
@@
To test the hypothesis 3 mapping of the mathematical framework (casual abstraction) to the model attention heads, various “interchange interventions” (aka activation patching) experiments were performed on the model, where
@@
Table 14: We define “positive-answer subtraction” and “negative-answer subtraction” calculation complexity quanta that parallel the addition quanta.
@@
We experimented with three approaches to re-using the trained addition model in the “mixed” (addition and subtraction) model:
@@
  • Freeze Attention: As per “Initialize Only”, but also every 100 training steps recopy the attention head weights from the addition model into the partially-trained mixed model.
  • Freeze All: As per “Initialize Only”, but also every 100 training steps recopy the entire addition model (attention heads and MLP layers) into the partially-trained mixed model.
Our intuition was that “Initialize Only” would give the mixed model the most freedom to learn new algorithms, but that the “Freeze Attention” and “Freeze All” approaches would make the resulting trained mixed model easier to interpret (as we could reuse our addition model insights).
After experimentation we found that the “Initialize Only” approach was the only one that quickly trained to be able to do both addition and subtraction accurately. We concluded that the other two methods constrain the model’s ability to learn new algorithms too much.
We also experimented with “where” in the model we inserted the addition (6-digit, 2-layer, 3-head) model into the slightly larger (6-digit, 3-layer, 4-head) mixed model. That is, do we initialize the first 2 layers or the last 2 layers of the mixed model? Also do we initialize the first 3 attention heads or the last 3 attention heads of the mixed model? Our intuition was that initializing the first layers and heads would be more likely to cause the model to re-use the addition circuits adding interpretability, so we used this approach.
@@
  • The model contains a new subtask that stands out: The algorithm relies on calculations done at token position P0, when the model has only seen one question token! What information can the model gather from just the first token? Intuitively, if the first token is a “8” or “9” then the first answer token is more likely to be a “+” (and not a “-”). The model uses this heuristic even though this probabilistic information is sometimes incorrect and so will work against the model achieving very low loss.